# DehazeGS: Seeing Through Fog with 3D Gaussian Splatting

### v4

@@ 4.1 Experimental Setup
Baselines. In addition to comparing with the original 3DGS Kerbl et al. (2023), we also compared our method with ScatterNeRF Ramazzina et al. (2023) and DehazeNeRF Chen et al. (2023). Furthermore, since SeaSplat Yang et al. (2024a) (a real-time rendering method for underwater scenes) also has some dehazing reconstruction capabilities, we included it in the comparison as well. Since DehazeNeRF has not released its source code, its qualitative and quantitative results are both sourced from its original paper. Note that the views of the real dataset presented in DehazeNeRF were specifically cropped (removing some edge regions of the scene). To ensure a fair comparison, we also cropped the edge regions of the rendered views for the novel perspectives accordingly.
@@
Training Details. All models, with the exception of DehazeNeRF, whose data is sourced from its original paper, all other models were trained on NVIDIA RTX 4090 GPU. For ScatterNeRF and SeaSplat, we followed the default settings provided in their respective papers. Due to the slower training speed of ScatterNeRF, we utilized three GPUs with a batch size of 2048 for training while keeping the total number of iterations consistent with the paper at 250K. SeaSplat was trained on a single GPU with the number of iterations set to 30K, consistent with the settings in its paper. For our model, training was performed on a single GPU. The number of iterations was set to 30K for synthetic datasets and 3K for real-world datasets. $\lambda_{D}$, $\lambda_{B}$, and $\lambda_{d_{rec}}$ were all set to 0.1. $\lambda_{d}$ utilized a continuous learning rate decay function adapted from JaxNeRF, where the initial weight was set to 1 at step 0, and the final weight was reduced to 0.01 at the specified maximum step, with logarithmically interpolated applied between. The learning rate for the convolutional network was set to $1\times 10^{-6}$, with a weight decay factor of 0.1, decaying over the corresponding number of iterations for synthetic and real scenes.
@@
Figure 3: Qualitative comparison of novel view synthesis results on real datasets, the images of DehazeNeRF are taken from its original paper. Our method exhibits finer texture details and is closer to the ground truth (GT) compared to existing methods. We encourage readers to zoom in to further explore the details further. In the supplementary materials, we present additional results of novel view synthesis.
@@
Dataset bear elephant lion Avg Time
Method—Metric PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
ScatterNeRF (250K) 9.53 0.302 0.754 9.8 0.339 0.786 8.21 0.326 0.772 $>$ 16 hours
SeaSplat (30K) 12.13 0.586 0.263 11.09 0.539 0.314 12.16 0.53 0.285 $\sim$ 25 mins
3DGS (30K) 13.09 0.597 0.276 13.28 0.605 0.307 12.01 0.507 0.337 $\sim$ 5 mins
DehazeNeRF (50K) 14.92 0.439 0.341 17.87 0.73 0.15 18.01 0.558 0.259 $>$ 4 hours
Ours (3K) 16.79 0.725 0.173 19.19 0.791 0.113 16.53 0.722 0.163 $\sim$ 1.2 mins
4.2 Quantitative and Qualitative Results
Table 1 and Table 2 present the quantitative comparison results on real hazy datasets and synthetic hazy datasets (The data in DehazeNeRF is sourced from its original paper. However, since it does not provide quantitative metrics for the bear and lion scenes, we opted to calculate them based on the rendered images presented in its paper). For real foggy datasets, Our method outperforms existing approaches in rendering details, achieving an average improvement of 40.4% in LPIPS, 29.7% in SSIM, and 3.3% in PSNR compared to the second-best method. Moreover, it achieves optimal performance within only 3K iterations, which requires approximately 1 minute, whereas NeRF-based methods typically take several hours. For the synthetic foggy dataset, our method outperforms the second-best method with improvements of 24.7% in LPIPS, 1.33% in SSIM, and 4.7% in PSNR. Additionally, our method achieves a rendering speed that is tens of times faster than that of NeRF-based methods.
Figure 4: Dehazing rendering results of the synthetic foggy dataset. Our method exhibits finer texture details and is closer to the ground truth (GT) compared to existing methods
For the qualitative comparison results, as shown in Fig. 3 and Fig. 4 (we encourage readers to zoom in to further explore the details further), the dehazing results of SeaSplat exhibit noticeable color anomalies, and are missing regions at the edges. Compared to DehazeNeRF, our reconstruction results outperform in both dehazing quality and the details of near and far scenes. Meanwhile, ScatterNeRF fails to reconstruct the scenes on both datasets. Additionally, our model is capable of achieving dehazing reconstruction in both indoor and outdoor complex scenarios.
@@ 4.2 Quantitative and Qualitative Results
Dataset garden bonsai counter stump Avg Time
Method—Metric PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
ScatterNeRF (250K) 14.15 0.346 0.609 12.57 0.394 0.689 12.46 0.385 0.751 11.82 0.224 0.771 $>$ 20 hours
3DGS (30K) 12.49 0.62 0.255 12.43 0.609 0.161 13.06 0.621 0.193 11.10 0.516 0.298 $\sim$ 10 mins
SeaSplat (30K) 18.42 0.687 0.275 20.4 0.705 0.121 19.16 0.736 0.183 17.78 0.631 0.293 $\sim$ 30 mins
Ours (30K) 18.72 0.742 0.183 22.19 0.726 0.104 19.31 0.704 0.153 19.10 0.637 0.219 $\sim$ 25 mins
Table 3: Ablation study on the synthetic foggy dataset. Our framework (DehazeGS) is employed starting from $\mathcal{L}_{rec}$.
@@
$\mathcal{L}_{DCP}+\mathcal{L}_{BCP}+\mathcal{L}_{rec}$ 19.74 0.697 0.175
$\mathcal{L}_{DCP}+\mathcal{L}_{BCP}+\mathcal{L}_{rec}+\mathcal{L}_{d}$ 19.77 0.692 0.166
@@
4.3 Ablation Study
In this section, we perform ablation studies on the various components of our framework to evaluate the role and contribution of each part. Specifically, we investigate the effects of depth regularization, depth-weighted, and relevant physical priors on the rendering results. The quantitative results are presented in Table 3. Note that Vanilla 3DGS refers to the version where the modeling of Gaussian distributions for foggy scenes is omitted from our framework (the top left section of Fig. 2), and only the reconstruction loss is used (The results of its novel view synthesis are shown as 3DGS in Fig. 3 and Fig. 4). Starting from $\mathcal{L}_{rec}$, it represents our framework, but with other loss functions removed, retaining only the reconstruction loss. We can see that when all the losses are combined, our total loss function achieves the maximum benefit. However, using only partial loss functions significantly reduces rendering quality.
5 Conclusion
We propose the first framework for learning clear 3D Gaussian splatting solely from multi-view hazy images. Our approach leverages explicit Gaussian representation to interpret the formation of hazy images through a physically accurate forward rendering process. This enables joint optimization of 3D scene representation while learning the participating medium. Overall, our method outperforms existing methods in terms of rendering quality, training speed, and rendering efficiency. However, there is still room for improvement. Although our method can reconstruct finer texture details at a faster speed and achieve better dehazing results, the dehazing of distant scenes still needs further enhancement in certain scenarios. We hope that our approach can inspire future work in related areas.

### v5

@@ Introduction
Baselines. In addition to comparing with the original 3DGS (Kerbl et al. 2023), we also compared our method with ScatterNeRF (Ramazzina et al. 2023) and DehazeNeRF (Chen et al. 2023). We further evaluated the reconstruction performance of combining low-level computer vision dehazing models with 3DGS. Specifically, we first dehazed the multi-view images using the pretrained models of Dehamer (Guo et al. 2022) or DehazeFormer (Song et al. 2023) (as shown in the supplementary materials), which were then fed into 3DGS for reconstruction. Furthermore, since SeaSplat (Yang, Leonard, and Girdhar 2024) (a real-time rendering method for underwater scenes) also has some dehazing reconstruction capabilities, we included it in the comparison as well. Since DehazeNeRF has not released its source code, its qualitative and quantitative results are both sourced from its original paper. Note that the views of the real dataset presented in DehazeNeRF were specifically cropped (removing some edge regions of the scene). To ensure a fair comparison, we also cropped the edge regions of the rendered views accordingly.
@@
Training Details. ScatterNeRF was trained on three 4090 GPUs with a batch size of 2048. All other models were trained on a single 4090 GPU, except for DehazeNeRF, whose results are taken from its original paper. SeaSplat was trained on a single GPU. All other parameters followed the default settings. For our model, training was performed on a single GPU. The number of iterations was set to 30K for synthetic datasets and 3K for real-world datasets. $\lambda_{D}$, $\lambda_{B}$, and $\lambda_{d_{rec}}$ were all set to 0.1. The learning rate for the convolutional network was set to $1\times 10^{-7}$, with a weight decay factor of 0.1, decaying over the corresponding number of iterations for synthetic and real scenes.
@@
Figure 3: Qualitative comparison of novel view synthesis (input foggy images and render clear images under novel views unseen by the model) results on real datasets, the images of DehazeNeRF are taken from its original paper. We encourage readers to refer to the supplementary materials to view rendering results from more views.
@@
Dataset bear elephant lion Avg PSNR Avg SSIM Avg LPIPS Avg Time
[…]
ScatterNeRF (250K) 10.34 0.351 0.728 11.01 0.431 0.727 9.92 0.320 0.763 10.42 0.367 0.739 >> 16 hours
SeaSplat (30K) 12.13 0.586 0.263 11.09 0.539 0.314 12.16 0.530 0.285 11.79 0.552 0.287 $\sim$ 25 mins
3DGS (30K) 13.09 0.597 0.276 13.28 0.605 0.307 12.01 0.507 0.337 12.79 0.570 0.307 $\sim$ 5 mins
Dehamer+3DGS (30K) 13.88 0.393 0.315 13.36 0.444 0.368 13.50 0.431 0.328 13.58 0.423 0.337 $\sim$ 5 mins
DehazeNeRF (50K) 14.92 0.439 0.341 17.87 0.730 0.150 18.01 0.558 0.259 16.93 0.576 0.250 >> 4 hours
Ours (3K) 16.79 0.725 0.173 19.19 0.791 0.113 17.33 0.709 0.186 17.77 0.742 0.157 $\sim$ 1.2 mins
Quantitative and Qualitative Results
Table 1 and Table 2 present the quantitative comparison results on real hazy datasets and synthetic hazy datasets. Since DehazeNeRF does not provide quantitative metrics for the bear and lion scenes, we calculate them based on the rendered images presented in its paper. On real foggy datasets, our method outperforms existing methods in rendering details, with average improvements of 37.2% (LPIPS), 28.8% (SSIM), and 5.0% (PSNR) over the second-best method. For the synthetic foggy dataset, it achieves 24.3% (LPIPS), 1.9% (SSIM), and 4.7% (PSNR) gains over the same baseline. Additionally, our method achieves a rendering speed that is tens of times faster than that of NeRF-based methods.
The qualitative comparison results are shown in Fig. 3 and Fig. 4. SeaSplat’s dehazing results show noticeable color anomalies and missing edge regions. Compared to DehazeNeRF, our reconstruction results outperform in both dehazing quality and the details of near and far scenes. Meanwhile, ScatterNeRF fails to reconstruct clear scenes on both datasets. Additionally, our model is capable of achieving dehazing reconstruction in both indoor and outdoor complex scenarios. For additional rendering results, please refer to Figures 1 and 2 in the supplementary materials. For the comparison results on the Waymo dataset, please refer to Figure 3 in the supplementary materials.
Figure 4: Qualitative comparison of dehazing results on the synthetic foggy dataset. Our method exhibits finer texture details and is closer to the ground truth (GT) compared to existing methods. In the supplementary materials, we provide other views and rendering results of the Waymo dataset.
@@
Dataset garden bonsai counter stump Avg PSNR Avg SSIM Avg LPIPS Avg Time
Method—Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS (30K) 12.49 0.620 0.255 12.43 0.609 0.161 13.06 0.621 0.193 11.10 0.516 0.298 12.27 0.592 0.227 $\sim$ 10 mins
ScatterNeRF (250K) 14.15 0.346 0.609 18.47 0.492 0.586 18.32 0.506 0.633 11.82 0.224 0.771 15.69 0.392 0.649 >> 20 hours
SeaSplat (30K) 18.42 0.687 0.275 20.40 0.705 0.121 19.16 0.736 0.183 17.78 0.631 0.293 18.94 0.689 0.218 $\sim$ 30 mins
Ours (30K) 18.72 0.742 0.183 22.19 0.726 0.104 19.31 0.704 0.153 19.10 0.637 0.219 19.83 0.702 0.165 $\sim$ 25 mins
Ablation Study
In this section, we perform ablation studies on the various components of our framework to evaluate the role and contribution of each part. Specifically, we investigate the effects of depth regularization ($\mathcal{L}_{d}$), depth-weighted ($\mathcal{L}_{d_{rec}}$), and relevant physical priors ($\mathcal{L}_{DCP}$, $\mathcal{L}_{BCP}$) on rendering results. The quantitative results are presented in Table 3. Note that Vanilla 3DGS (Kerbl et al. 2023) refers to the original and unmodified 3DGS, without the ASM model. (the results of its novel view synthesis are shown as 3DGS in Fig. 3 and Fig. 4). Starting from $\mathcal{L}_{rec}$, which corresponds to our framework with only the reconstruction loss retained and all other loss terms removed, we observe that the performance progressively improves as additional losses are incorporated.
Figure 5: Qualitative ablation results show progressive improvements from vanilla GS to the complete model, with enhanced details and dehazing in distant and nearby scenes.
Table 3: Ablation study on the synthetic foggy dataset. Each component of our framework (DehazeGS) is employed starting from $\mathcal{L}_{rec}$.
@@
$\mathcal{L}_{rec}+\mathcal{L}_{DCP}+\mathcal{L}_{BCP}$ 19.74 0.697 0.175
$\mathcal{L}_{rec}+\mathcal{L}_{DCP}+\mathcal{L}_{BCP}+\mathcal{L}_{d}$ 19.77 0.692 0.166
@@
Conclusion
We propose the first framework for learning clear 3D Gaussian splatting solely from multi-view hazy images. Our approach leverages explicit Gaussian representation, based on the atmospheric scattering model, to interpret the formation of hazy images through a physically accurate forward-rendering process. This enables joint optimization of 3D scene representation while learning the participating medium. Overall, our method outperforms existing methods in terms of rendering quality, training speed, and rendering efficiency. However, in certain scenarios, the dehazing effect on distant scenes still needs further improvement. We hope that our approach can inspire future research in this field.
Acknowledgements
This research was supported by the National Natural Science Foundation of China (62202076), the National Key R&D Program of China under Grant 2022ZD0160804, National Natural Science Foundation of China (U21A20515, 52238003, U22B2034, 82174224), Shenzhen S&T programme (No. CJGJZD20240729141906008).